OpenStudy (anonymous):

the ratio of females to males on a particular flight was 2 : 3. Females represented five more than 1/3 of all people aboard. how many people were on that flight

OpenStudy (anonymous):

x=number of women, y=number of men. 2x=3y x-5=1/3(x+y)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I would set females equal to "x" and males equal to "y" and the total number of people on the plane equal to "z" Remember that our target variable is z. which gives us: 2x = 3y and x = 1/3(z) + 5 and x + y = z (this says that the total number of people aboard is the sum of males and females) Solve by substitution.